the triangle with sides of 5,8,9, is acute, right or obtuse
Answer by khwang(438)

Among 5,8 & 9 , the largest is 9 side. Let the angle opposite to the side 9
is A which is also the largest angle . By the cosine law:
9^2 = 5^2 + 8^2 - 2*5*8 cos A = 25 + 64 - 80 cos A = 89 - 80 cos A
So, cos A = 8/80 =1/10 > 0.
Hence, A is acute and the triangle is acute.